OpenStudy (anonymous):

A recipe asks for 2 1/3 cup of flour. How many cups of flour are needed for 3 1/2 recipes?

OpenStudy (inowalst):

\[2\frac{ 1 }{ 3 } \times 3\frac{ 1 }{ 2 } = x\] Multiply 2 times 3. Then 1 times 1. Then 2 times 3.

OpenStudy (anonymous):

6 1/6?

OpenStudy (inowalst):

Yes.

OpenStudy (anonymous):

my teacher is checking them stay here and if its right ill give u a medal

OpenStudy (anonymous):

61/1

OpenStudy (inowalst):

61/1 ? What?